# Set and functional prediction: randomness, exchangeability, and conformal

### v1

@@ Abstract
This paper continues the study of the efficiency of conformal prediction as compared with more general randomness prediction and exchangeability prediction. It does not restrict itself to the case of classification, and our results will also be applicable to the case of regression. The price to pay is that efficiency will be attained only on average, albeit with respect to a wide range of probability measures on the label space.
@@ 1 Introduction
Conformal prediction is usually presented as a method of set prediction [10, Part I], i.e., as a way of producing prediction sets (rather than point predictions). Another way to look at a conformal predictor is as a way of producing a p-value function (discussed, in a slightly different context, in, e.g., [4]), which is a function mapping each possible label $y$ of a test object to the corresponding conformal p-value. In analogy with “prediction sets”, we will call such p-value functions “prediction functions”. The prediction set $\Gamma^{\alpha}$ corresponding to a prediction function $f$ and a significance level $\alpha\in(0,1)$ (our target probability of error) is the set of all labels $y$ such that $f(y)>\alpha$. A standard property of validity for conformal predictors is that $\Gamma^{\alpha}$ makes an error (fails to cover the true label) with probability at most $\alpha$; it is implied by the conformal p-values being bona fide p-values (under suitable assumptions, such as data exchangeability).
The most standard assumption in machine learning is that of randomness (i.e., the data are assumed to be produced in the IID fashion). This paper is a development of [8], which introduces the most general class of predictors, “randomness predictors”, that produce prediction functions that are valid, in the same sense as conformal predictors, under the assumption of randomness. There are many more randomness predictors than conformal predictors, and an interesting question is whether there are randomness predictors that have significant advantages (e.g., in practice) over conformal predictors. This question was answered (albeit imperfectly) in [8] in the case of classification with few (such as two) classes. In this paper we will be interested in arbitrary label spaces, including the case of regression. The message of this paper is similar to that of [8]: the difference between conformal and randomness prediction is not huge, but it remains an open question whether it can be usefully exploited. This paper strengthens most of the positive results in [8], but its only negative result is much weaker (but also simpler) than the negative results of [8].
@@
A useful technical tool in establishing connections between conformal and randomness predictors is provided by conformal e-predictors [9], which are obtained by replacing p-values with e-values. Conformal e-predictors output e-value functions $f$ as their prediction functions. Such functions $f$ can also be represented in terms of the corresponding prediction sets $\Gamma^{\alpha}:=\{y\mid f(y)<\alpha\}$, where $\alpha\in(0,\infty)$ is the significance level (notice that now we exclude the labels with large e-values from the prediction set, which is opposite to what we did for p-values). However, the property of validity of conformal e-predictors is slightly more difficult to state in terms of prediction sets: now validity means that the integral of the probability of error for $\Gamma^{\alpha}$ over $\alpha\in(0,\infty)$ does not exceed 1 [9, end of Appendix B]. This implies that the probability of error for $\Gamma^{\alpha}$ is at most $1/\alpha$, but this simple derivative property of validity is much weaker.
@@
Conformal e-predictors are not only a useful technical tool, but we can also use them for prediction directly. In Glenn Shafer’s opinion [7], e-values are even more intuitive than p-values. Because of the importance of e-predictors, in the rest of this paper we will use the word “predictor” in combinations such as “conformal predictor” and “randomness predictor” generically, including both p-predictors (standard predictors based on p-values) and e-predictors (predictors based on e-values); in particular, we will never drop “p-” in “p-predictor”.
In this paper we will follow the scheme for establishing the closeness of conformal and randomness predictors used in [8, Figure 1]. Namely, we will establish connections between:
  1. (a)
    conformal p-predictors and conformal e-predictors;
  2. (b)
    conformal e-predictors and exchangeability e-predictors, where the class of exchangeability e-predictors is intermediate between the conformal e-predictors and the randomness e-predictors, and it consists of randomness e-predictors that are valid under the assumption of exchangeability (which is weaker than the assumption of randomness in the case of finite data sequences);
  3. (c)
    exchangeability e-predictors and randomness e-predictors;
  4. (d)
    randomness e-predictors and randomness p-predictors.
Steps (a) and (d) (converting p-values to e-values and back) are standard, and we will mainly concentrate on steps (b) and (c).
We start in Sect. 2 from the main definitions, and Sect. 3 is devoted to the main results. In particular, we establish the efficiency of conformal predictors among randomness predictors in both p- and e-versions. Namely, the prediction functions for conformal predictors turn out to be competitive on average with the prediction functions for any randomness predictors, where “on average” refers to an arbitrary probability measure that can depend on the test example. Sections 4 and 5 give some applications to classification and regression, respectively, and Sect. 6 concludes.
@@ 2 Definitions
This paper is about the following prediction problem (continuing the discussion started in [8]). We are given a training sequence of examples $z_{i}=(x_{i},y_{i})$, $i=1,\dots,n$ for a fixed $n$, each consisting of an object $x_{i}$ and its label $y_{i}$, and a new test object $x_{n+1}$; the task is to predict $x_{n+1}$’s label $y_{n+1}$. A potential label $y$ for $x_{n+1}$ is true if $y=y_{n+1}$ and false otherwise. The objects are drawn from a non-empty measurable space $\mathbf{X}$, the object space, and the labels from the label space $\mathbf{Y}$, which is assumed to be a non-trivial measurable space (meaning that the $\sigma$-algebra on it is different from $\{\emptyset,\mathbf{Y}\}$).
A measurable function $P:\mathbf{Z}^{n+1}\to[0,1]$ is a randomness p-variable if, for any probability measure $Q$ on $\mathbf{Z}$ and any significance level $\alpha\in(0,1)$, $Q^{n+1}(\{P\leq\alpha\})\leq\alpha$. And a measurable $P:\mathbf{Z}^{n+1}\to[0,1]$ is a conformal p-variable if
  • $R(\{P\leq\alpha\})\leq\alpha$ for any exchangeable probability measure $R$ on $\mathbf{Z}^{n+1}$ and any $\alpha\in(0,1)$;
  • it is training-invariant, i.e., invariant w.r. to permutations of the training examples:
    $P(z_{\sigma(1)},\dots,z_{\sigma(n)},z_{n+1})=P(z_{1},\dots,z_{n},z_{n+1})$ (1)
    for each data sequence $z_{1},\dots,z_{n+1}$ and each permutation $\sigma$ of $\{1,\dots,n\}$ (training-invariant functions were called simply invariant in [8]).
We will sometimes refer to the values taken by p-variables as p-values, and our notation for the classes of all randomness and conformal p-variables will be $\mathcal{P}^{\mathrm{R}}$ and $\mathcal{P}^{\mathrm{tX}}$, respectively.
Conformal p-variables can be used for prediction, and we will also refer to them as conformal p-predictors. Notice that the standard expression “training set” is only justified for predictors $P$ satisfying (1) (and even in this case it is not justified completely; it would be more accurate to say “training bag”). There are several ways to package the output of conformal p-predictors. One is in terms of set prediction: for each significance level $\alpha\in(0,1)$, each training sequence $z_{1},\dots,z_{n}$, and each test object $x_{n+1}$, we can output the prediction set
@@
By the definition of a conformal p-variable, under the assumption of exchangeability, the probability that a conformal p-predictor makes an error at significance level $\alpha$, i.e., the probability of $y_{n+1}\notin\Gamma^{\alpha}$, is at most $\alpha$. See [8, Sect. 2] for a more detailed discussion of connections between conformal p-variables and conformal p-prediction.
Instead of predicting with one prediction set in the family (2), we can package our prediction as the prediction function
@@
We may refer to this mode of prediction as functional prediction. The step from set prediction to functional prediction is analogous from the step from confidence intervals to p-value functions (see, e.g., [6, Sect. 9] and [3, 4, 5] for the latter).
@@
Similarly, we can use randomness p-variables for prediction, and then we refer to them as randomness p-predictors. By definition, the probability that the prediction set (2) derived from a randomness p-predictor makes an error is at most $\alpha$, this time under the assumption of randomness. We will use the prediction functions (3) for randomness p-predictors as well.
Two important desiderata for conformal and randomness predictors are their validity and efficiency. In terms of the prediction function $f$, validity concerns the value $f(y_{n+1})$ at the true label (its typical values should not be too small in p-prediction), and efficiency concerns the values $f(y)$ at the false labels $y\neq y_{n+1}$ (they should be as small as possible in p-prediction). Validity is automatic under randomness (and even exchangeability for conformal predictors), and in this paper we are interested in the efficiency of conformal predictors relative to other randomness predictors. Later in the paper (Theorems 7 and 8 below) we will establish efficiency guarantees for conformal prediction in terms of randomness prediction.
A nonnegative measurable function $E:\mathbf{Z}^{n+1}\to[0,\infty]$ is a randomness e-variable if $\int E\,\mathrm{d}Q^{n+1}\leq 1$ for any probability measure $Q$ on $\mathbf{Z}$. It is an exchangeability e-variable if $\int E\,\mathrm{d}R\leq 1$ for any exchangeable probability measure $R$ on $\mathbf{Z}^{n+1}$. We will denote the classes of all randomness and exchangeability e-variables by $\mathcal{E}^{\mathrm{R}}$ and $\mathcal{E}^{\mathrm{X}}$, respectively. The class of all measurable functions $E:\mathbf{Z}\to[0,\infty$ is denoted by $\mathcal{E}$.
The class $\mathcal{E}^{\mathrm{tX}}$ of conformal e-variables consists of all functions $E\in\mathcal{E}^{\mathrm{X}}$ that are training-invariant:
$E(z_{\sigma(1)},\dots,z_{\sigma(n)},z_{n+1})=E(z_{1},\dots,z_{n},z_{n+1})$ (4)
for each data sequence $z_{1},\dots,z_{n+1}$ and each permutation $\sigma$ of $\{1,\dots,n\}$. We often regard the randomness e-variables $E\in\mathcal{E}^{\mathrm{R}}$ as randomness e-predictors and conformal e-variables $E\in\mathcal{E}^{\mathrm{tX}}$ as conformal e-predictors. Similarly to (3), they output prediction functions
@@
For conformal and randomness e-predictors, validity and efficiency change direction: for validity, typical values $f(y_{n+1})$ should not be too large, and for efficiency typical values $f(y)$ at the false labels $y\neq y_{n+1}$ should be as large as possible. Again validity is automatic under randomness, and Theorem 7 below establishes efficiency guarantees for conformal e-prediction in terms of randomness e-prediction.
We will also need two important subclasses of $\mathcal{E}^{\mathrm{R}}$. The subclass $\mathcal{E}^{\mathrm{tR}}$ of $\mathcal{E}^{\mathrm{R}}$ consists of all functions $E\in\mathcal{E}^{\mathrm{R}}$ that are training-invariant (i.e., satisfy (4)). The subclass $\mathcal{E}^{\mathrm{iR}}$ of $\mathcal{E}^{\mathrm{R}}$ consists of all functions $E\in\mathcal{E}^{\mathrm{R}}$ that are invariant w.r. to all permutations:
@@
for each permutation $\pi$ of $\{1,\dots,n+1\}$; let us call such randomness e-variables invariant (this is almost the same thing as configuration randomness e-variables in [8]).
A big advantage of e-variables over p-variables is that the average of e-variables is again an e-variable. This allows us to define, given an e-variable $E\in\mathcal{E}^{\mathrm{R}}$, three derivative e-variables:
@@
$\pi$ ranging over the permutations of $\{1,\dots,n+1\}$ and $\sigma$ ranging over the permutations of $\{1,\dots,n\}$. It is clear that $E^{\mathrm{i}}\in\mathcal{E}^{\mathrm{iR}}$ whenever $E\in\mathcal{E}^{\mathrm{R}}$, that $E^{\mathrm{X}}\in\mathcal{E}^{\mathrm{X}}$ for all $E\in\mathcal{E}$, and that $E^{\mathrm{t}}\in\mathcal{E}^{\mathrm{tX}}$ whenever $E\in\mathcal{E}^{\mathrm{X}}$. The operators (5) and (7) are kinds of averaging: while $E\mapsto E^{\mathrm{i}}$ averages over all permutations of an input data sequence (including both training and test examples), $E\mapsto E^{\mathrm{t}}$ averages over the permutations of the training sequence only.
Using two of these three operators, we can turn any randomness e-variable $E$ to an exchangeability e-variable $E^{\mathrm{X}}$ to a conformal e-variable $(E^{\mathrm{X}})^{\mathrm{t}}$. The following lemma shows that the order in which the last two operators are applied does not matter.
Lemma 2.
Proof.
Let us fix a data sequence $z_{1},\dots,z_{n+1}$ and check $(E^{\mathrm{t}})^{\mathrm{X}}(z_{1},\dots,z_{n+1})=(E^{\mathrm{X}})^{\mathrm{t}}(z_{1},\dots,z_{n+1})$. As functions of a permutation of $z_{1},\dots,z_{n+1}$, $E$ and $E^{\mathrm{X}}$ are proportional to each other, and therefore, $E^{\mathrm{t}}$ and $(E^{\mathrm{X}})^{\mathrm{t}}$ are also proportional to each other. This implies $(E^{\mathrm{t}})^{\mathrm{X}}=(E^{\mathrm{X}})^{\mathrm{t}}$ on the permutations of $z_{1},\dots,z_{n+1}$. And this is true for each $(z_{1},\dots,z_{n+1})$. ∎
We will let tX stand for the composition of the two operators:
$E^{\mathrm{tX}}:=(E^{\mathrm{t}})^{\mathrm{X}}=(E^{\mathrm{X}})^{\mathrm{t}}.$
It is easy to see that $E\in\mathcal{E}$ belongs to $\mathcal{E}^{\mathrm{X}}$ if and only if, for any data sequence $z_{1},\dots,z_{n+1}$,
$\frac{1}{(n+1)!}\sum_{\pi}E(z_{\pi(1)},\dots,z_{\pi(n+1)})\leq 1,$ (8)
$\pi$ ranging over the permutations of $\{1,\dots,n+1\}$. Let us say that such an $E$ is admissible if (8) always holds with “$=$” in place of $\leq$”. (This agrees with the standard notion of admissibility in statistical decision theory.)
The intuition (which can be formalized easily) behind the operators that we have just introduced is that:
  • t projects $\mathcal{E}^{\mathrm{X}}$ onto $\mathcal{E}^{\mathrm{tX}}$;
In particular, these operators are idempotent:
$(E^{\mathrm{i}})^{\mathrm{i}}=E^{\mathrm{i}},\quad(E^{\mathrm{X}})^{\mathrm{X}}=E^{\mathrm{X}},\quad(E^{\mathrm{t}})^{\mathrm{t}}=E^{\mathrm{t}},\quad(E^{\mathrm{tX}})^{\mathrm{tX}}=E^{\mathrm{tX}}.$
Despite these operators being projections, we cannot claim that these ways of moving between different function classes are always optimal.

### v2

# wideformula

@@ Abstract
Conformal predictors provide set or functional predictions that are valid under the assumption of randomness, i.e., under the assumption of independent and identically distributed data. The question asked in this paper is whether there are predictors that are valid in the same sense under the assumption of randomness and that are more efficient than conformal predictors. The answer is that the class of conformal predictors is universal in that only limited gains in predictive efficiency are possible. The previous work in this area has relied on the algorithmic theory of randomness and so involved unspecified constants, whereas this paper’s results are much more practical. They are also shown to be optimal in some respects.
@@ 1 Introduction
The main assumption of machine learning is that of randomness, i.e., the assumption that the observations are independent and identically distributed (IID). The method of conformal prediction [1, 17] allows us to complement predictions output by standard machine-learning algorithms by some measures of confidence in those predictions that are valid under the assumption of randomness. However, conformal prediction only uses the assumption of exchangeability, which is weaker than randomness. A natural question is how limiting this is; does conformal prediction lose much by not using the full strength of the assumption of randomness and only using exchangeability? The main message of this paper and the earlier paper by Nouretdinov et al. [9] is that the answer is “no” (we don’t lose much), at least in the first approximation. In the terminology of [9], conformal prediction is universal.
The main limitation of the pioneering paper [9] is that it is based on the algorithmic theory of randomness. Because of that, Nouretdinov et al.’s results involve unspecified constants and, therefore, are never applicable to practical machine learning. This paper’s results are more precise in several respects, but the main one is that they only involve fully specified constants and, therefore, open up the possibility of quantifying limitations of conformal predictors. Other advances of this paper as compared with [9] are that our results are not restricted to the case of classification, but in the case of classification they are stronger and are complemented by optimality results.
To demonstrate that conformal prediction does not lose much by not using the full strength of the assumption of randomness, this paper introduces the most general class of predictors, “randomness predictors”, which produce predictions of the same kind as conformal predictors that are only required to be valid under the assumption of randomness. There are many more randomness predictors than conformal predictors, and our question is whether conformal predictors are as good as arbitrary randomness predictors. The answer given in this paper is a qualified yes: every randomness predictor $P$ can modelled by a conformal predictor $P^{\prime}$ so that the predictions output by $P^{\prime}$ are almost as good as those output by $P$. It is unclear whether the difference between conformal and randomness prediction can be usefully exploited at all [16].
One simplifying assumption made in this paper is that it concentrates on predictors that are train-invariant, i.e., invariant w.r. to permutations of the training sequence. (We will remove this assumption only in Appendix B.) The assumption of train-invariance for predictors is extremely natural under the assumption of randomness for the data. It is reflected in the standard expression for a training sequence being “training set” in machine learning; since “set” implies the lack of order, this expression is only justified for train-invariant predictors (and even in this case it is not justified completely; it would have been more accurate to say “training bag”). In fact, conformal predictors can be defined as train-invariant predictors that are valid under exchangeability [9, Proposition 1]. This will be discussed in detail in Sect. 2, where we will also see that the requirement of train-invariance is justified by general principles of statistical inference.
@@
Conformal prediction is usually presented as a method of set prediction [17, Part I], i.e., as a way of producing prediction sets (rather than point predictions). Another way to look at a conformal predictor is as a way of producing a p-value function (discussed, in a slightly different context, in, e.g., [5]), which is a function mapping each possible label $y$ of a test object to the corresponding conformal p-value. In analogy with “prediction sets”, we may call such p-value functions “prediction functions”. The prediction set $\Gamma^{\alpha}$ corresponding to a prediction function $f$ and a significance level $\alpha\in(0,1)$ (our target probability of error) is the set of all labels $y$ such that $f(y)>\alpha$. A standard property of validity for conformal predictors is that $\Gamma^{\alpha}$ makes an error (fails to cover the true label) with probability at most $\alpha$; it is implied by the conformal p-values being bona fide p-values (under suitable assumptions, such as data exchangeability).
@@
To establish connections between conformal and randomness predictors we will use conformal e-predictors [15], which are obtained by replacing p-values with e-values (for the definition of e-values, see, e.g., [6, 10, 18] or Sect. 2 below). Conformal e-predictors output e-value functions $f$ as their prediction functions. Such functions $f$ can also be represented in terms of the corresponding prediction sets $\Gamma^{\alpha}:=\{y\mid f(y)<\alpha\}$, where $\alpha\in(0,\infty)$ is the significance level (notice that now we exclude the labels with large e-values from the prediction set, which is opposite to what we did for p-values). However, the property of validity of conformal e-predictors is slightly more difficult to state in terms of prediction sets: now validity means that the integral of the probability of error for $\Gamma^{\alpha}$ over $\alpha\in(0,\infty)$ does not exceed 1 [15, end of Appendix B]. This implies that the probability of error for $\Gamma^{\alpha}$ is at most $1/\alpha$, but this simple derivative property of validity is much weaker.
@@
Conformal e-predictors are not only a useful technical tool, but we can also use them for prediction directly. In Shafer’s opinion [11], e-values are even more intuitive than p-values. Because of the importance of e-predictors, in the rest of this paper we will use the word “predictor” in combinations such as “conformal predictor” and “randomness predictor” generically, including both p-predictors (standard predictors based on p-values) and e-predictors (predictors based on e-values); in particular, we will never drop “p-” in “p-predictor”. This is a potential source of confusion, and the reader should keep in mind that the usual notion of conformal predictors corresponds to conformal p-predictors in this paper.
$\mathcal{E}^{\mathrm{tR}}$$\mathcal{P}^{\mathrm{tR}}$$\mathcal{P}^{\mathrm{tX}}$$\mathcal{E}^{\mathrm{tX}}$
Figure 1: A square representing the main classes of predictors considered in this paper: $\mathcal{E}^{\mathrm{tR}}$ (train-invariant randomness e-predictors), $\mathcal{E}^{\mathrm{tX}}$ (conformal e-predictors), $\mathcal{P}^{\mathrm{tR}}$ (train-invariant randomness p-predictors), $\mathcal{P}^{\mathrm{tX}}$ (conformal p-predictors).
We start in Sect. 2 from the main definitions, including those of conformal and randomness predictors. The main classes of predictors that we are interested in in this paper are shown in Figure 1. All four classes are train-invariant. An arrow going from class $A$ to class $B$ means embedding: $A\subseteq B$. Section 3 is devoted to the main results, first for e-predictors in Subsect. 3.1 and then for p-predictors in Subsect. 3.2. A line between classes $A$ and $B$ in Figure 1 means the possibility of transformations from $f\in A$ to $f^{\prime}\in B$ and vice versa; such transformations, called “calibration”, will be discussed at the beginning of Subsect. 3.2 and will serve as a way of deducing results for p-predictors from those for e-predictors. In the two subsections, we establish the predictive efficiency of conformal predictors among randomness predictors in both e- and p-versions. Namely, the prediction functions for conformal predictors turn out to be competitive on average with the prediction functions for any randomness predictors, where “on average” refers to an arbitrary probability measure that can depend on the test example. These results are illustrated on the simple case of binary classification. Section 4 gives applications to multi-class classification, and Sect. 5 concludes listing some limitations of our results and directions of further research.
Our notation for the base of natural logarithms will be $\mathrm{e}\approx 2.72$ (while italic $e$ will often serve as a generic notation for e-values).
@@ 2 Definitions
This paper deals with the following prediction problem. We are given a training sequence of examples $z_{i}=(x_{i},y_{i})$, $i=1,\dots,n$ for a fixed $n$, each consisting of an object $x_{i}$ and its label $y_{i}$, and a new test object $x_{n+1}$; the task is to predict $x_{n+1}$’s label $y_{n+1}$. A potential label $y$ for $x_{n+1}$ is true if $y=y_{n+1}$ and false otherwise. The objects are drawn from a non-empty measurable space $\mathbf{X}$, the object space, and the labels from the label space $\mathbf{Y}$, which is assumed to be a non-trivial measurable space (meaning that the $\sigma$-algebra on it is different from $\{\emptyset,\mathbf{Y}\}$).
A measurable function $P:\mathbf{Z}^{n+1}\to[0,1]$ is a randomness p-variable if, for any probability measure $Q$ on $\mathbf{Z}$ and any significance level $\alpha\in(0,1)$, $Q^{n+1}(\{P\leq\alpha\})\leq\alpha$. Such a function $P$ is an exchangeability p-variable if $R(\{P\leq\alpha\})\leq\alpha$ for any exchangeable probability measure $R$ on $\mathbf{Z}^{n+1}$ and any $\alpha\in(0,1)$. And a real-valued function $P$ defined on $\mathbf{Z}^{n+1}$ is train-invariant if it is invariant w.r. to permutations of the training examples:
$P(z_{\sigma(1)},\dots,z_{\sigma(n)},z_{n+1})=P(z_{1},\dots,z_{n},z_{n+1})$
for each data sequence $(z_{1},\dots,z_{n+1})\in\mathbf{Z}^{n+1}$ and each permutation $\sigma$ of $\{1,\dots,n\}$. In other words, train-invariant functions should depend on the training examples $z_{1},\dots,z_{n}$ only via the training bag $\lbag z_{1},\dots,z_{n}\rbag$. Finally, $P$ is a conformal p-variable if it is a train-invariant exchangeability p-variable.
Remark 1.
The difference between the assumptions of randomness and exchangeability disappears for infinite data sequences under mild assumptions about the example space $\mathbf{Z}$ (it is required to be Borel). This follows from de Finetti’s theorem, which represents exchangeable probability measures as integral mixtures of product probability measures $Q^{\infty}$. The difference becomes essential for finite data sequences; see, e.g., [17, Sections 2.9.1 and A.5.1].
We will sometimes refer to the values taken by p-variables as p-values, and our notation for the classes of all randomness, train-invariant randomness, exchangeability, and conformal p-variables will be $\mathcal{P}^{\mathrm{R}}$, $\mathcal{P}^{\mathrm{tR}}$, $\mathcal{P}^{\mathrm{X}}$, and $\mathcal{P}^{\mathrm{tX}}$, respectively.
Conformal p-variables can be used for prediction, and we will also refer to them as conformal p-predictors (they are usually called simply “conformal predictors” [1, 17]). There are several ways to package the output of conformal p-predictors, as discussed in Sect. 1. One is in terms of set prediction: for each significance level $\alpha\in(0,1)$, each training sequence $z_{1},\dots,z_{n}$, and each test object $x_{n+1}$, we can output the prediction set
@@
By the definition of a conformal p-variable, under the assumption of exchangeability, the probability that a conformal p-predictor makes an error at significance level $\alpha$, i.e., the probability of $y_{n+1}\notin\Gamma^{\alpha}$, is at most $\alpha$.
Instead of predicting with one prediction set in the family (2), in this paper we prefer to package our prediction as the prediction function
@@
We may refer to this mode of prediction as functional prediction. The step from set prediction to functional prediction is analogous to the step from confidence intervals to p-value functions (see, e.g., [8, Sect. 9] and [4, 5] for the latter).
Remark 2.
There are several equivalent definitions of conformal p-predictors, and the definition as train-invariant exchangeability p-variables (first given in [9, Proposition 1]) is one of them. Let us check that it is equivalent to, e.g., the definition of conformal p-values given in [17, (2.20)] for a fixed length $n$ of the training sequence. In one direction, it is obvious that conformal p-values are train-invariant and valid under exchangeability. On the other hand, given a train-invariant exchangeability p-variable $P$, we can define the nonconformity measure
$A(\lbag z_{1},\dots,z_{n+1}\rbag,z_{i}):=1/P(z_{i+1},\dots,z_{n+1},z_{1},\dots,z_{i}),$
and the resulting conformal p-values will dominate $P$ (dominate in the sense of being less than or equal to).
@@
Similarly, we can use randomness p-variables for prediction, and then we refer to them as randomness p-predictors. By definition, the probability that the prediction set (2) derived from a randomness p-predictor makes an error is at most $\alpha$, this time under the assumption of randomness. We will use the prediction functions (3) for randomness p-predictors as well. Less important, we call exchangeability p-variables exchangeability p-predictors.
Two standard desiderata for conformal, and by extension randomness, predictors are their validity and efficiency. In terms of the prediction function $f$, validity concerns the value $f(y_{n+1})$ of $f$ at the true label (the typical values should not be too small in p-prediction), and efficiency concerns the values $f(y)$ at the false labels $y\neq y_{n+1}$ (they should be as small as possible in p-prediction). Validity is automatic under randomness (and even under exchangeability for conformal predictors), and in this paper we are interested in the efficiency of conformal predictors relative to other randomness predictors. Later in the paper (Theorem 5 and Corollary 7 below) we will establish efficiency guarantees for conformal prediction in terms of randomness prediction.
A nonnegative measurable function $E:\mathbf{Z}^{n+1}\to[0,\infty]$ is a randomness e-variable if $\int E\,\mathrm{d}Q^{n+1}\leq 1$ for any probability measure $Q$ on $\mathbf{Z}$. It is an exchangeability e-variable if $\int E\,\mathrm{d}R\leq 1$ for any exchangeable probability measure $R$ on $\mathbf{Z}^{n+1}$. We will denote the classes of all randomness and exchangeability e-variables by $\mathcal{E}^{\mathrm{R}}$ and $\mathcal{E}^{\mathrm{X}}$, respectively. The class of all measurable functions $E:\mathbf{Z}^{n+1}\to[0,\infty]$ is denoted by $\mathcal{E}$. It is easy to see that $E\in\mathcal{E}$ belongs to $\mathcal{E}^{\mathrm{X}}$ if and only if, for any data sequence $z_{1},\dots,z_{n+1}$,
$\frac{1}{(n+1)!}\sum_{\pi}E(z_{\pi(1)},\dots,z_{\pi(n+1)})\leq 1,$ (4)
$\pi$ ranging over the permutations of $\{1,\dots,n+1\}$.
The class $\mathcal{E}^{\mathrm{tX}}$ of conformal e-variables consists of all functions $E\in\mathcal{E}^{\mathrm{X}}$ that are train-invariant. We often regard the randomness e-variables $E\in\mathcal{E}^{\mathrm{R}}$ as randomness e-predictors and conformal e-variables $E\in\mathcal{E}^{\mathrm{tX}}$ as conformal e-predictors. Similarly to (3), they output prediction functions
@@
Remark 4.
Similarly to Remark 2, it is easy to check that a train-invariant exchangeability e-variable is the same thing as a conformal e-predictor as defined in, e.g., [15]. Indeed, the nonconformity e-measure [15, Sect. 2] corresponding to a train-invariant exchangeability e-variable $E$ is {wideformula*} A(z_1,…,z_n+1) := ( E(z_2,…,z_n+1,z_1), E(z_3,…,z_n+1,z_1,z_2), …,
E(z_1,z_2,…,z_n+1) ).
The subclass $\mathcal{E}^{\mathrm{tR}}\subseteq\mathcal{E}^{\mathrm{R}}$ of all train-invariant randomness e-predictors is important since under the assumption of randomness it is natural to consider only train-invariant predictors: the requirement of train-invariance is a special case of the principle of sufficiency in statistical inference [2, 2.3.(ii)]. Let us refer to the requirement of train-invariance under the assumption of randomness as the train-invariance principle. The train-invariance principle is a special case not only of the sufficiency principle but also of the invariance principle [2, Example 2.35], which makes it even more convincing.
For conformal and randomness e-predictors, validity and efficiency change direction as compared with p-predictors: for validity, typical values $f(y_{n+1})$ should not be too large, and for efficiency typical values $f(y)$ at the false labels $y\neq y_{n+1}$ should be as large as possible. Again validity is automatic under randomness, and Theorem 5 below establishes efficiency guarantees for conformal e-prediction in terms of train-invariant randomness e-prediction. Similarly to $\mathcal{E}^{\mathrm{R}}$ and $\mathcal{E}^{\mathrm{tX}}$, elements of the class $\mathcal{E}^{\mathrm{X}}$ will be referred to as exchangeability e-predictors, but we are not particularly interested in them per se.
We will also need another, even narrower, subclass of $\mathcal{E}^{\mathrm{R}}$, $\mathcal{E}^{\mathrm{iR}}\subseteq\mathcal{E}^{\mathrm{tR}}$. The class $\mathcal{E}^{\mathrm{iR}}$ consists of all e-variables $E\in\mathcal{E}^{\mathrm{R}}$ that are invariant w.r. to all permutations:
@@
for each permutation $\pi$ of $\{1,\dots,n+1\}$; let us call such randomness e-variables invariant. An equivalent definition is to say that $E(z_{1},\dots,z_{n+1})$ depends on the data sequence $z_{1},\dots,z_{n+1}$ only via the bag $\lbag z_{1},\dots,z_{n+1}\rbag$ of its elements. The interpretation is that the value $E(z_{1},\dots,z_{n+1})$ for $E\in\mathcal{E}^{\mathrm{iR}}$ is the degree to which we can reject the hypothesis that the bag $\lbag z_{1},\dots,z_{n+1}\rbag$ resulted from its elements being generated in the IID fashion.
A big advantage of e-variables over p-variables is that the average of e-variables is again an e-variable. This allows us to define, given an e-variable $E\in\mathcal{E}^{\mathrm{R}}$, two derivative e-variables:
@@
with $\pi$ ranging over the permutations of $\{1,\dots,n+1\}$ and $/0$ 0 interpreted as 1 (the last convention may be relevant to (6)). More generally, we can allow any $E\in\mathcal{E}$ in (5) and (6). It is clear that $E^{\mathrm{i}}\in\mathcal{E}^{\mathrm{iR}}$ whenever $E\in\mathcal{E}^{\mathrm{R}}$ and that $E^{\mathrm{X}}\in\mathcal{E}^{\mathrm{X}}$ for all $E\in\mathcal{E}$. For $E\in\mathcal{E}^{\mathrm{R}}$, the operator (5) performs averaging of e-values over all permutations of an input data sequence, and (6) is the relative deviation of an e-value from the average.
In the most important for us case (transforming train-invariant randomness e-predictors to conformal e-predictors), the operator (6) is polynomially computable: namely, if $E\in\mathcal{E}^{\mathrm{tR}}$ is efficiently computable, the extra time when computing $E^{\mathrm{X}}\in\mathcal{E}^{\mathrm{tX}}$ on top of computing $E$ is linear, $O(n)$. This follows from
$E^{\mathrm{X}}(z_{1},\dots,z_{n+1})=\frac{E(z_{1},\dots,z_{n+1})}{\frac{1}{n+1}\sum_{i=1}^{n+1}E(z_{i+1},\dots,z_{n+1},z_{1},\dots,z_{i})}$
for $E\in\mathcal{E}^{\mathrm{tR}}$.